\begin{thebibliography}{00}
\bibitem{ACB} E. Abdalla and L. Correa-Borbonet, The elliptic solutions to the Friedmann equation and the Verlinde's maps, (2002), e-print  arXiv:hep-th/0212205.

\bibitem{AS} R. Aurich and F. Steiner, The cosmic microwave background for a nearly flat compact hyperbolic universe, {\it Monthly Notices Royal Astron. Soc.} {\bf 323} (2001), 1016-1024.

\bibitem{AST} R. Aurich, F. Steiner, and H. Then, Numerical computation of Mass waveforms and an application to cosmology, from Proceedings of the International School on Mathematical Aspects of Quantum Chaos II, Lecture Notes in Physics, Springer-Verlag, Berlin (2004), e-print  arXiv:gr-qc/0404020.

\bibitem{BBDR} G. Bag, B. Bhui, S. Das, and F. Rahaman, A study on Bianchi IX cosmological model in Lyra geometry, {\it Fizika} {\bf B 12} (2003), 193-200.

\bibitem{BCHKR} N. Begum, S. Charkraborty, M. Hossain, M. Kalam, and F. Rahaman, Bianchi IX string cosmological model in Lyra geometry, {\it Pramana J. of Physics} {\bf 60} (2003), 1153-1159.

\bibitem{Biermann} G. Biermann, Probelemata quaedam mechanica functionum ellipticarum ope soluta, Dissertatio Inauguralis, Friedrich Wilhelm Universit\"at (1865).
%, available at http://edoc.hu-berlin.de/ebind/hdiss/BIERPROB\_PPN313151385/XML.

\bibitem{BF} P. Byrd and M. Friedman, Handbook of Elliptic Integrals for Engineers and Physicists, {\it Grundlehren der mathematischen Wissenschaften} {\bf 67}  Springer-Verlag, Berlin  (1954).

\bibitem{Chandra} K. Chandrasekharan, Elliptic Functions, {\it Grundlehren der mathematischen Wissenschaften} {\bf 281} (1985), Springer-Verlag, Berlin.

\bibitem{JDMSRI} J. D'Ambroise, Applications of elliptic and theta functions to Friedmann-Robertson-Lema\^itre-Walker Cosmology, from A Window into Zeta and Modular Physics, Editors K. Kirsten and F. Williams, MSRI Pub. {\bf 57} Cambridge University Press (2010), 279-293, e-print arXiv:gr-qc/0908.2481.


\bibitem{JDthesis} J. D'Ambroise, Generalized EMP and nonlinear Schr\"odinger-type reformulations of some scalar field cosmological models, Ph.D. Thesis, Univ. of Massachusetts at Amherst (2010), e-print arXiv:gr-qc/1005.1410.

\bibitem{JDFW} J. DÕAmbroise and F. L. Williams, A dynamic correspondence between Bose-Einstein condensates and FLRW and Bianchi I cosmology with a cosmological constant, {\it Journal of Mathematical Physics} {\bf 51} (2010), No. 6, 062501-062511, e-print arXiv:math-ph/1007.4237 .


\bibitem{Greenhill} A. Greenhill, The Applications of Elliptic Functions, Dover Publications (1959).


\bibitem{Kharb} L. Kharbediya, Some exact solutions of the Friedmann equations with the cosmological term, {\it Astronom. Zh.} {\bf 53} (1976), 1145-1152. 

\bibitem{KT} G.Konopelchenko and I.Taimanov , Generalized Weierstrass formulae, soliton equations and Willmore surfaces : I .Tori of revolution and the mKDV equation, (1995), e-print arXiv :dg-ga/9506011.

\bibitem{KW} G. Kraniotis and S. Whitehouse, General relativity, the cosmological constant and modular forms, {\it Classical and Quantum Gravity} {\bf 19} (2002) 5073-5100, e-print arXiv:gr-qc/0105022.

\bibitem{InhomCosMod} A. Krasi\'nski, Inhomogeneous Cosmological Models, Cambridge Univ. Press (1997).

\bibitem{Kras} A. Krasi\'nski, Physics and cosmology in an inhomogeneous universe, from Proceedings of the $49^{th}$ Yamada Conference on Black Holes and High-Energy Astrophysics, Universal Academy Press, Tokyo (1998), 133-147, e-print arXiv:gr-qc/9806039.

\bibitem{Lemaitre} G. Lema\^itre, The Expanding Universe, {\it Ann. Soc. Sci. Bruxelles} {\bf A 53}, 81 (1933).  English translation:  {\it Gen. Relativity and Gravitation} {\bf 29} (1997), 641-680.


\bibitem{LidseyBEC} Lidsey, J., Cosmic Dynamics of Bose-Einstein Condensates, {\it{ Classical and Quantum Gravity}} {\bf 21} (2004), 777-785, e-print arXiv:gr-qc$\backslash$0307037.

\bibitem{Omer} G. Omer, Jr., Spherically symmetric distribution of matter without pressure, {\it Proc. National Acad. of Sciences} {\bf 53} (1965), 1-5.

\bibitem{Reynolds} M. Reynolds, An exact solution in non-linear oscillations, {\it J.Physics A:Math. Gen.} {\bf 22} (Letter to the Editor) (1989), L723-L726.

\bibitem{WW}E. Whittaker and G. Watson, A Course of Modern Analysis, Cambridge Mathematical Library, Cambridge University Press (1927).


\end{thebibliography}

\documentclass{ws-jnmp}

\usepackage{amssymb, amsmath, textcomp, dsfont, mathrsfs, verbatim}

\begin{document}


\markboth{Jennie D'Ambroise and Floyd L. Williams}
{Parametric solution of certain nonlinear differential equations in cosmology}

%
%
\copyrightauthor{J. D'Ambroise and F. L. Williams}

\title{PARAMETRIC SOLUTION OF CERTAIN NONLINEAR DIFFERENTIAL EQUATIONS IN COSMOLOGY}

\author{\footnotesize JENNIE D'AMBROISE}

\address{Division of Science and Mathematics, University of Minnesota - Morris\\
Morris, Minnesota 56267, USA\\
\email{jdambroi@morris.umn.edu}}

\author{FLOYD L. WILLIAMS}

\address{Department of Mathematics and Statistics, University of Massachusetts\\
Amherst, Massachusetts 01003, USA\\
\email{williams@math.umass.edu}}

\maketitle

\begin{history}
\received{(Day Month Year)}
\revised{(Day Month Year)}
\accepted{(Day Month Year)}
\end{history}

\begin{abstract}
We obtain in terms of the Weierstrass elliptic $\wp-$function, sigma function, and zeta function an explicit parametrized solution of a particular nonlinear, ordinary differential equation.  This equation includes, in special cases, equations that occur in the study of both homogeneous and inhomogeneous cosmological models, and also in the dynamic Bose-Einstein condensates - cosmology correspondence, for example. 
\end{abstract}

\keywords{Weierstrass P-function; Bianchi cosmological models; elliptic functions.}

\ccode{2000 Mathematics Subject Classification: 33E05, 83C20, 83F05}

\section{Introduction}	

In this paper we solve the nonlinear differential equation
\begin{equation}\dot{Y}(t)^2=\frac{f(Y(t))}{Y(t)^{2n}},\label{eq: Yeqngeneraln}\end{equation}
where $f(x)=a_0x^4+4a_1x^3+6a_2x^2+4a_3x+a_4$ is a quartic polynomial with no repeated factors and $n\geq 0$ is a fixed whole number.  The solution is expressed parametrically in terms of the Weierstrass $\wp$-function $\wp(w)$, and his sigma and zeta functions $\sigma(w), \zeta(w)$; see formulas (2.4), (2.5), (2.6), (2.7), and definitions (A.1), (A.5) of the Appendix.  One can also solve equation (1.1) in case $f(x)$ does have a repeated factor -- a situation  which is easier to deal with and which, in particular, generally does not involve elliptic functions; see Example 6.

Of some special interest is the equation
\begin{equation}\dot{Y}(t)^2=BY(t)^2+EY(t)-K+\frac{A}{Y(t)}+\frac{D}{Y(t)^2},\label{eq: Yeqn}\end{equation}
with $n=1$ in (1.1).  Note that the solution of the differential equation
\begin{equation}\dot{Y}_1(t)^2=Y_1(t)^2\left[ B_1+E_1 Y_1(t)^m - K_1 Y_1(t)^{2m}+A_1 Y_1(t)^{3m}+D_1 Y_1(t)^{4m}\right],\label{eq: Yeqnwithm}\end{equation}
where $m\neq 0$ is a fixed whole number, can be obtained  from that of (\ref{eq:  Yeqn}).  Namely, if we set $Y(t)=Y_1(t)^{-m}$ then equation (\ref{eq: Yeqnwithm}) is transformed to (\ref{eq: Yeqn}) for $B=m^2B_1, E=m^2E_1, K=m^2K_1, A=m^2A_1$, and $D=m^2D_1$.  

If both $E$ and $D$ are zero, for example, then the general solution $Y_{E,D=0}$ of (\ref{eq: Yeqn}), in terms of the Weierstrass elliptic function $\wp(w)$, was obtained in 1933 by G. Lema\^itre \cite{Lemaitre}, in his study of spherically symmetric distributions of matter.  Compare also the paper \cite{Omer} of G. Omer with references therein to special case solutions by R. Tolman, B. Datta, and H. Bondi.  The solution $Y_{E,D=0}$ provides for an exact, inhomogeneous cosmological solution $ds^2$ of the Einstein field equations with cosmological constant $\Lambda=3B$.  Namely, for the family of \emph{Szekeres-Szafron solutions}

\begin{equation}ds^2=dt^2-e^{2B(x,y,z,t)}(dx^2+dy^2)-e^{2A(x,y,z,t)}dz^2,\label{eq: SSmetric}\end{equation}
the functions $B(x,y,z,t), A(x,y,z,t)$ are explicated by the solution $Y_{E,D=0}$.  The reference \cite{KW}, for example, contains a detailed discussion of this matter, and the text \cite{InhomCosMod} can be consulted for a comprehensive analysis of inhomogeneous cosmology.  Also see \cite{Kras}, where one can consider $D=-Q^2\neq 0$ (with yet $E=0$), $Q$ being a constant electric charge.

The Friedmann-Lema\^itre-Robertson-Walker (FLRW) metric can be obtained from $ds^2$ in (\ref{eq: SSmetric}) (as a ``limit"), and a known formula (see formula (42) of \cite{KW}, for example) for the scale factor (the ``radius" of the FLRW universe) also follows from the general formulas presented here.  Further remarks on this, as well as the computation of scale factors in anisotropic models (in the Bianchi $V$ and $IX$ models, for example) are taken up in section 3.

Another case of interest, among others to be mentioned later, is that when both $A$ and $B$ are zero in (\ref{eq: Yeqn}).  Here the solution $Y_{A,B=0}$ is of relevance regarding the dynamic correspondence between Bose-Einstein condensates (BECs) and FLRW/Bianchi $I$ cosmology \cite{LidseyBEC, JDFW}.  In particular we deduce an alternate formula (in section 3) for the second moment $I_2(t)=Y_{A,B=0}(t)$ of the wavefunction of the Gross-Pitaevskii equation, for BECs governed by a time-dependent, harmonic trapping potential - especially when a cosmological constant is present - i.e. when $E\neq 0$.  The second moment in fact determines the harmonic trapping frequency.

Our formulas therefore provide for a general, unifying context where some known formulas in the literature are immediately derived or extended, and some new ones are developed -- as seen in the concrete examples of section 3.

\section{The Solution Of Equation (\ref{eq: Yeqngeneraln})}

\indent\indent For  $Z\stackrel{def.}{=} \sqrt{ a_0 Y^4+4a_1Y^3+6a_2Y^2+4a_3Y+a_4}$ equation (1.1) is expressed as 
\begin{equation}t=\displaystyle\int\frac{Y^n}{Z}dY+\delta\end{equation}
for an integration constant $\delta$.   By a choice of any root $x_0$ of $f(x_0)=0$ and a (finite) Taylor expansion $f(x)=4\alpha_3(x-x_0)+6\alpha_2(x-x_0)^2+4\alpha_1(x-x_0)^3+\alpha_0(x-x_0)^4$ of $f(x)$ about $x_0$, where $\alpha_3=f'(x_0)/4, \alpha_2=f''(x_0)/12$, $\alpha_1=f'''(x_0)/24$, $\alpha_0=f''''(x_0)/24$ one can reduce the elliptic integral $I$ in (2.1) to a \emph{Weierstrass canonical form}, where the quartic in $Z$ is reduced to a cubic.  Namely, by the substitution $Y=x_0+\alpha_3/(x-\alpha_2/2)$ one gets 
\begin{equation}I = \displaystyle\int \frac{\left( x_0+\frac{\alpha_3}{x-\alpha_2/2}\right)^n}{\sqrt{4x^3-g_2x-g_3}} dx\end{equation}
where
\begin{eqnarray}
g_2&=& 3\alpha_2^2-4\alpha_1\alpha_3=a_0a_4-4a_1a_3+3a_2^2,\\
g_3&=&2\alpha_1\alpha_2\alpha_3-\alpha_2^3-\alpha_0\alpha_3^2=a_0a_2a_4+2a_1a_2a_3-a_2^3-a_0a_3^2-a_1^2a_4\notag\end{eqnarray}
are the \emph{Weierstrass invariants} of $f(x)$.  Here $\alpha_3\neq 0$ since $f'(x)\neq 0$, as $f(x)$ has no repeated factors, by hypothesis.  Moreover if $\wp(w)=\wp(w; g_2, g_3)$ is the \emph{Weierstrass $\wp$-function} attached to $g_2, g_3$ (as in the Appendix) then, by equation (A.2) there, the substitution $x=\wp(w+c)$, for any fixed constant $c$, leads to the re-statement
\begin{equation}t=\displaystyle\int \left( x_0 + \frac{f'(x_0)}{4\left[ \wp(w+c)-f''(x_0)/24\right]}\right)^n dw+\delta\end{equation}
of equation (2.1), where we also have (by the above substitution $Y\rightarrow x$)
\begin{equation} Y=x_0+\frac{f'(x_0)}{4\left[ \wp(w+c)-f''(x_0)/24\right]}.\end{equation}

Equations (2.4), (2.5), which constitute the main result, provide for a parametric solution of equation (1.1).  For the applications we have in mind, in examples 1-6 below, we need only the cases $n=0, 1, 2$.  For $n=0$ and the choice $\delta=0$, $t=w$ in (2.4) and formula (2.5) also follows from a general formula of Biermann-Weierstrass \cite{Biermann}; more on this point later.  Let $\sigma(w), \zeta(w)$ denote the Weierstrass \emph{sigma} and \emph{zeta functions} respectively; see definition (A.5).  For a choice $w_0$ with $\wp(w_0)=f''(x_0)/24$ we have for $n=1, 2$ \emph{respectively} in (2.4)
\begin{equation}t=x_0w + \frac{f'(x_0)}{4\wp'(w_0)}\left[ \log \frac{\sigma(w+c-w_0)}{\sigma(w+c+w_0)} + 2(w+c)\zeta(w_0) \right] +\delta, \label{eq: n=1integration} \end{equation}
\begin{eqnarray}t&=&x_0^2w+\left[ - \frac{x_0 f'(x_0)}{2\wp'(w_0)}+\frac{f'(x_0)^2\wp''(w_0)}{16 \wp'(w_0)^3}\right] \log \frac{\sigma(w+c+w_0)}{\sigma(w+c-w_0)}\notag\\
&& - \frac{f'(x_0)^2}{16\wp'(w_0)^2}\left[ \zeta(w+c+w_0)+\zeta(w+c-w_0)\right]\label{eq: n=2integration}\\
&& + (w+c)\left( \frac{x_0 f'(x_0)}{\wp'(w_0)}\zeta(w_0)-\frac{f'(x_0)^2}{16}\left[ \frac{2\wp(w_0)}{\wp'(w_0)^2}+\frac{2\wp''(w_0)\zeta(w_0)}{\wp'(w_0)^3}\right] \right) + \delta,\notag
\end{eqnarray}
by formulas 1037.06, 1037.11, respectively, in \cite{BF} -- assuming also that $\wp(w_0)\neq$ the roots $e_1, e_2, e_3$ of $4x^3-g_2x-g_3=0$.  In fact, in principle, one can compute the integral in (2.4) for \emph{arbitrary} $n\geq 0$, using formulas in \cite{BF}, which we shall not pursue as the result is not needed here.

The aforementioned Biermann-Weierstrass formula for the solution $Y(t)$ of equation (1.1) in case $n=0$ is
\begin{equation}Y(t)=Y(0)+ \frac{  \left[ f(Y(0))^{1/2}\wp'(t)+\frac{f'(Y(0))}{2}\left( \wp(t)-\frac{f''(Y(0))}{24}\right)+\frac{f(Y(0))f'''(Y(0))}{24}\right]}{2\left[ \wp(t)-\frac{f''(Y(0))}{24}\right]^2-\frac{f(Y(0))f''''(Y(0))}{48}}.\end{equation}
Also see references \cite{Reynolds, WW}.

\section{Some examples}

The following examples are meant to provide application and further clarity of the preceding formulas.

\begin{example}
We begin with equation (20)
\begin{equation}\dot\Phi(t)^2=-K(z)+\frac{2M(z)}{\Phi(t)}+\frac{\Lambda \Phi(t)^2}{3}\label{eq: SSeqn}\end{equation}
of \cite{KW}, which is associated with the metric (\ref{eq: SSmetric}) in the introduction.  Since the functions $K(z), M(z)$ are independent of $t$, equation (\ref{eq: SSeqn}) is mathematically the same as equation (8.9) of \cite{Lemaitre}, and since its solution is presented and  discussed in \cite{KW} we greatly limit our remarks here.  (\ref{eq: SSeqn}) is equation (\ref{eq: Yeqn}) with $3B=\Lambda$ (a cosmological constant), $E=0, K=K(z), A=2M(z), D=0$.  Thus $f(x)=\left( \Lambda/3\right) x^4 -K(z)x^2+2M(z)x$ has $x_0=0$ as a first order root for $M(z)\neq 0$, since $f'(0)=2M(z)$.  By formulas (2.5), (2.6) the solution of (\ref{eq: SSeqn}) is given parametrically by
\begin{eqnarray}
\Phi&=& \frac{M(z)/2}{\wp(w+c; g_2, g_3)+K(z)/12}\notag\\
&&\label{eq: SSparamsoln}\\
t&=&\frac{M(z)}{2\wp'(w_0)}\left[ \log \frac{\sigma(w+c-w_0)}{\sigma(w+c+w_0)}+2(w+c)\zeta(w_0)\right]+\delta\notag\end{eqnarray}
for $\wp(w_0)=-K(z)/12\neq e_1, e_2, e_3$, which are formulas (21), (23) of \cite{KW}, where $g_2=K(z)^2/12, g_3=K(z)^3/216-\Lambda M(z)^2/12$ (by (2.3)),  and where now the integration constant $\delta=\delta(z)$ depends on the variable $z$.

\end{example}

\begin{example}
Consider the Friedmann equation
\begin{equation}a'(\eta)^2=-\kappa a(\eta)^2+\frac{K A_r}{c^4}+\frac{KA_m}{c^3}a(\eta)+\frac{c^2\Lambda}{3}a(\eta)^4\label{eq: EFEconformaltime}\end{equation}
for the scale factor $a(\eta)$ in a FLRW universe whose energy and matter are modeled by a perfect fluid.  Here $\eta$ is \emph{conformal time}, $\kappa$ is the curvature parameter, $K=8\pi G/3$ for the Newton constant $G$, $A_r$ and $A_m$ are radiation and matter constants, $c$ is the speed of light, and $\Lambda$ is a cosmological constant.  Given the ``big bang" initial condition $a(0)=0$ one derives immediately from (2.8) the result
\begin{equation}a(\eta)=\frac{\frac{\left(KA_r\right)^{1/2}}{c^2}\wp'(\eta)+\frac{KA_m}{2c^3}\left( \wp(\eta)+\kappa/12\right)}{2\left(\wp(\eta)+\kappa/12\right)^2-K\Lambda A_r/6c^2},\label{eq: aetasolution}\end{equation}
with known application in the study of cosmic microwave background fluctuation, for example; see \cite{AS, AST}.  The invariants $g_2, g_3$ in (2.3) are given by 
\begin{eqnarray}
g_2&=& \frac{\Lambda K A_r}{3c^2}+\frac{\kappa^2}{12},\notag\\
&&\\
g_3&=&-\frac{\kappa \Lambda K A_r}{18c^2}+\frac{\kappa^3}{216}-\frac{K^2A_m^2\Lambda}{48c^4}.\notag\end{eqnarray}
Elliptic function solutions of Friedmann equations are also discussed in \cite{JDMSRI, Kharb, ACB}, for example.
\end{example}

\begin{example}  As another example we consider the Bianchi V cosmological model with metric
\begin{equation}ds^2=-dt^2+X(t)^2dx^2+e^{2b x}Y(t)^2dy^2+e^{2b x}Z(t)^2dz^2\end{equation}
for $b\neq 0$.  As before, we take the energy momentum tensor to be that of a perfect fluid and denote the radiation and matter constants by $A_r, A_m$, respectively.  For $K=8\pi G/3$, a zero cosmological constant $\Lambda=0$, and the speed of light taken to be $c=1$, the Einstein equation is a special case of (1.1)
\begin{equation}\dot{R}(t)^2=\frac{1}{R(t)^4}\left[b^2 R(t)^4+KA_m R(t)^3+KA_r R(t)^2+KD\right]\end{equation}
in terms of $R(t)\stackrel{def.}{=}\left(X(t)Y(t)Z(t)\right)^{1/3}$ and the quantity $D\stackrel{def.}{=}$ $\frac{R(t)^2}{9K}\left(\frac{\dot{X}^2}{X^2}+\frac{\dot{Y}^2}{Y^2}+\frac{\dot{Z}^2}{Z^2}-\frac{\dot{X}\dot{Y}}{XY}\right.$ $\left.-\frac{\dot{X}\dot{Z}}{XZ}-\frac{\dot{Y}\dot{Z}}{YZ}\right)$ which can be shown to be a constant; see \cite{JDthesis}.    Then by (2.5) and (2.7) we obtain 
\begin{eqnarray}
R&=&x_0+\frac{f'(x_0)}{4\left[\wp(w+c)-f''(x_0)/24\right]}\nonumber\\
&&\\
t&=&x_0^2 w+\left( - \frac{x_0 f'(x_0)}{2\wp'(w_0)}+\frac{f'(x_0)^2\wp''(w_0)}{16\wp'(w_0)^3}\right) \log \frac{\sigma(w+c+w_0)}{\sigma(w+c-w_0)}\notag\\
&& -\frac{f'(x_0)^2}{16\wp'(w_0)^2}\left[ \zeta(w+c+w_0)+\zeta(w+c-w_0)\right] \notag\\
&&+(w+c)\left( \frac{x_0f'(x_0)}{\wp'(w_0)}\zeta(w_0)-\frac{f'(x_0)^2}{16}\left[\frac{2\wp(w_0)}{\wp'(w_0)^2}+\frac{2\wp''(w_0)\zeta(w_0)}{\wp'(w_0)^3}\right]\right)+\delta\notag
\end{eqnarray}
for $\wp(w_0)\stackrel{def.}{=}f''(x_0)/24$ and associated polynomial $f(x)=b^2 x^4+KA_m x^3+KA_r x^2+KD$.  The invariants are
\begin{eqnarray}
g_2&=&K\left(b^2D+\frac{A_r^2 K}{12}\right) \notag\\
&&\\
\notag g_3&=&\frac{K^2}{6}\left(b^2A_r D-\frac{A_r^3K}{36}-\frac{3DA_m^2K}{8}  \right),\end{eqnarray}
where as in the Appendix we assume that $g_2^3-27g_3^2\neq 0$.  To obtain the metric one writes $X=R, Y=Re^{\sqrt{3DK}\tau}, Z=Re^{-\sqrt{3DK}\tau}$ where $\tau(t)=\int dt/{R(t)^3}$.  That is, $\tau$ is given parametrically in terms of $w$ by
\begin{equation}\tau=
\frac{w}{x_0}-\frac{f'(x_0)}{4x_0^2\wp'(w_1)}\left[ \log\frac{\sigma(w+c-w_1)}{\sigma(w+c+w_1)}+2(w+c)\zeta(w_1)\right] + \delta' 
\end{equation}
for integration constant $\delta'$, $\wp(w_1)\stackrel{def.}{=}f''(x_0)/2-f'(x_0)/4x_0$ and assuming $x_0\neq 0$.

\end{example}

\begin{example}
For a cosmological model with Bianchi IX metric
\begin{equation}ds^2=-dt^2+a(t)^2dx^2+b(t)^2dy^2+\left[ b(t)^2\sin^2 y+a(t)^2\cos^2 y\right]dz^2-2a(t)^2\cos y \  dxdz,\label{eq: BIXmetric}\end{equation}
and massless scalar field $\Phi$ and flat (constant) potential $V(r)=2\lambda$, the modified Einstein equations based on Lyra geometry are studied in \cite{BBDR}, for example.  Also compare the paper \cite{BCHKR}.  The field equations yield the relation $\dot\Phi=\Phi_0/ab^2$ for an integration constant $\Phi_0$, and the assumption $a=b^n$ leads to the Einstein equation 
\begin{equation}\frac{\ddot{b}}{b}+(n+1)\frac{\dot{b}^2}{b^2}=\frac{1}{(n-1)b^2}-\frac{b^{2n-4}}{n-1}\label{eq: BIXbeqn}\end{equation}
for $n\neq 1$, which moreover is shown to have the first integral 
\begin{equation}\dot{b}^2=\frac{1}{n^2-1}-\frac{b^{2n-2}}{2n^2-2n}+D_1b^{-2n-2}\label{eq: BIXbeqnintegrated}\end{equation}
for $n\neq 0, \pm 1$, where $D_1$ is an integration constant.  The authors found solutions of (\ref{eq: BIXbeqnintegrated}) only for $D_1=0$ - for $n=2, 1/2, 3/2, 3/4$.  Therefore we consider the case $D_1\neq 0$, and we take $n=2$ for example:  $\dot{b}^2=1/3-b^2/4+D_1/b^6$, or $a=b^2 \Rightarrow$
\begin{equation}\dot{a}(t)^2=\frac{4}{3}a(t)-a(t)^2+\frac{D}{a(t)^2},\label{eq: BIXaeqn}\end{equation}
which is another example of equation (\ref{eq: Yeqn}) for $D=4D_1, A=0, K=0, E=\frac{4}{3}, B=-1.$  Here
\begin{eqnarray}
f(x)&=&-x^4+\frac{4}{3}x^3+4D_1,\label{eq: BIXinvars}\\
g_2&=&-4D_1, g_3=-\frac{4D_1}{9}=\frac{g_2}{9}.\notag\end{eqnarray}
In particular $g_2^3-27g_3^2=-16D_1^2(4D_1+\frac{1}{3})\neq 0$ for $D_1\neq 0, -\frac{1}{12}$, which we assume.  Certainly if $D_1=-\frac{1}{12}$, then $f(x)=-x^4+\frac{4}{3}x^3-\frac{1}{3}=-(x-1)^2\left(x^2+\frac{2}{3}x+\frac{1}{3}\right)$ has $x_0=1$ as a repeated root, for example.

On the other hand, consider $x_0=-1$ which is a non-repeated root for $D_1=\frac{7}{12}: \ f'(x_0)=8\neq 0$.  By (2.5), (2.6)   we have the parametric solution
\begin{eqnarray}
a&=&-1+\frac{2}{\wp(w+c)+5/6},\notag\\
&&\\
t&=&-w+\frac{2}{\wp'(w_0)}\left[ \log \frac{\sigma(w+c-w_0)}{\sigma(w+c+w_0)}+2(w+c)\zeta(w_0)\right]+\delta\notag\end{eqnarray}
of equation (\ref{eq: BIXaeqn}) for $\wp(w_0)=-5/6$.  One can find solutions to equation (\ref{eq: BIXaeqn}) for many other non-zero values of $D_1$ similarly.  This amounts to specifying $D_1\neq -\frac{1}{12}$ and solving for a corresponding root $x_0$ of $f(x)$ in (\ref{eq: BIXinvars}).
\end{example}

\begin{example}
The differential equation
\begin{equation}\dot{Y}(t)^2=EY(t)-K+\frac{D}{Y(t)^2},\label{eq: BECYeqn}\end{equation}
which is another example of equation (\ref{eq: Yeqngeneraln}) (for $n=1$), arises in the study of time-dependent, harmonically trapped Bose-Einstein condensates, as indicated in the introduction.  The constant $E$ here is a positive multiple of a $d-$dimensional cosmological constant, for $d\geq 3$ arbitrary.  Also $D$ is positive and $E\neq 0$.  See equation (35) in \cite{JDFW}, where some elliptic function solutions are discussed.  
Equations (2.5), (2.6) provide for a parametric solution of (\ref{eq: BECYeqn}).  However, as $f(x)=Ex^3-Kx^2+D$ is \emph{cubic} in this case, one has an alternate, simpler parametrization which in particular does not involve a logarithm, as in (2.6).

Namely consider the simple substitution $Y=ax+b$, $a\neq 0$, a suggestion for which we thank one of the referees.  Equation (2.1) then reads
\begin{eqnarray}
t&=& \displaystyle\int \frac{(a^2 x+ab)dx}{\sqrt{Ea^3x^3+(3Ea^2b-Ka^2)x^2+(3Eab^2-2Kab)x+Eb^3-Kb^2+D}}+\delta\nonumber\\
&&\\
&=& \frac{1}{a^2}\displaystyle\int \frac{(u+ab)du}{\sqrt{4u^3-g_2 u-g_3}}+\delta,\nonumber\end{eqnarray}
for $u=a^2x, a=(E/4)^{1/3}, b=K/3E, g_2'=(-3Eb^2+2Kb)/a$, and $g_3'=-Eb^3+Kb^2-D$.  Note that $g_2', g_3'$ are \emph{not} the invariants $g_2=K^2/12, g_3=\frac{K^3}{216}-\frac{E^2D}{16}$ of $f(x)$ given in definition (2.3).  Similar to the derivation of equation (2.4), we now let $u=\wp(w)=\wp(w; g_2', g_3')$ to get the parametric solution
\begin{eqnarray}
t &=& -\frac{1}{a^2}\zeta(w)+\frac{b}{a}w+\delta\nonumber\\
&&\\
Y&=&\frac{1}{a}\wp(w)+b\nonumber\end{eqnarray}
of equation (\ref{eq: BECYeqn}), where again $\zeta(w)$ is the Weierstrass zeta function of (A.5).

\end{example}

\begin{example}
In the main result, equations (2.4), (2.5), $f(x)$ is assumed to have no repeated factors.  However, if repeated factors occur then equation (1.1) can be solved, in fact, in terms of elementary, non-elliptic functions, which we illustrate in the following example.

The equation 
\begin{equation}U'(x)^2+4U(x)^4-2U(x)^2-U(x)/\sqrt{2}=\frac{1}{16},\end{equation}
an example of equation (1.1) with $n=0$, is satisfied by the potential $U(x)$ of a particular \emph{Zakharov-Shabat system}.  $U(x)$, moreover, satisfies a type of static, \emph{modified Novikov-Veselov} (mNV) \emph{equation} \cite{KT}, a point which we return to later.  Here $f(x)=-4x^4+2x^2+x/\sqrt{2}+\frac{1}{16}$ has $(x-x_0)$ as a repeated factor for $x_0\stackrel{def.}{=}-\sqrt{2}/4$, $f(x)=(x-x_0)^2\left[ B_1+B_2(x-x_0)+B_3(x-x_0)^2\right]$ being its finite Taylor expansion about $x_0$ for $2B_1=f''(x_0), 6B_2=f'''(x_0), 24 B_3=f''''(x_0)$.  By the substitution $x=x_0+\frac{1}{u}$,
\begin{equation} I\stackrel{def.}{=}\displaystyle\int \frac{dx}{\sqrt{f(x)}}=\pm \displaystyle\int \frac{du}{\sqrt{B_1 u^2+B_2 u+B_3}}\end{equation}
is an elementary integral (compare the integral in (2.1)) whose evaluation depends on the signs of $B_1$ and the discriminant $\Delta\stackrel{def.}{=}B_2^2-4B_1B_3$.  Actually, as $x_0=-\sqrt{2}/4, B_1=-1<0$ and $\Delta = 16 >0$. 

 In the end, for 
\begin{equation}I = -\frac{1}{\sqrt{-B_1}}arcsin\left( \frac{B_2+2B_1 u}{\sqrt{\Delta}}\right)\end{equation}
and an integration constant $\delta$ we obtain the solution
\begin{equation}U(x)=U_{\mp}(x; \delta) = \frac{\mp sin(x-\delta)}{ 2\sqrt{2}\left[ \sqrt{2}\pm sin(x-\delta)\right]  }\end{equation}
of equation (3.20).  In particular, one has the solution $U_+(x; 0)$ obtained in \cite{KT} (by a quite different method) where among other results the authors there demonstrate invariance of the \emph{Willmore functional} $W$ (the Polyakov extrinsic string action) under NV deformations. 

We note that the ``deformation"
\begin{equation}U(x,t)\stackrel{def.}{=}U_+(x+2t; 0) \stackrel{def.}{=} \frac{sin(x+2t)}{2\sqrt{2}\left[ \sqrt{2} - sin(x+2t) \right]}\end{equation}
of $U_+(x; 0)$ is a solution of the mNV equation
\begin{equation}U_t=U_{xxx}+24U^2 U_x.\end{equation}
The functional $W$ of course is a basic quantity in the study of two-dimensional gravity.

\end{example}




\appendix

\section{}

Given the central importance of the Weierstrass $\wp-$function $\wp(w)$ for the present work we recall briefly, for the reader's convenience, its construction/definition.  As we have indicated in section 2 a more detailed account is available in \cite{Chandra, Greenhill, WW}.  

Let $\omega_1, \omega_2$ be non-zero complex numbers.  Since the imaginary parts of a non-zero complex number $z$ and its reciprocal are related by $Im \ z^{-1}=-\left( Im \ z\right) \left| z \right|^{-2}$, one has that $Im \ \omega_2/\omega_1\neq 0$ if and only if $Im \ \omega_1/\omega_2\neq 0$.  In particular we assume that $Im \ \omega_2/\omega_1>0$, which is equivalent to the assumption $Im \ \omega_1/\omega_2<0$.  The corresponding \emph{lattice} $\mathscr{L}=\mathscr{L}(\omega_1, \omega_2)$ generated by $\omega_1$ and $\omega_2$ is defined to be the set of points $\omega=m\omega_1+n\omega_2$ where $m$ and $n$ vary over the set of whole numbers.  The lattice $\mathscr{L}$ gives rise to the $\wp-$function
\begin{equation}\wp(w)\stackrel{def.}{=}\frac{1}{w^2}+\displaystyle\sum_{\omega\in\mathscr{L}-\{0\}}\left[ \frac{1}{(w-\omega)^2}-\frac{1}{\omega^2}\right]\label{eq: defnwp}\end{equation}
which is also denoted by $\wp(w; \mathscr{L})$, or by $\wp(w; \omega_1, \omega_2)$.  $\wp(w)$ is a meromorphic function, which is doubly periodic with periods $\omega_1, \omega_2$.  Thus, by definition, $\wp(w)$ is an \emph{elliptic} function.  $\wp(w)$ has double poles at $w=\omega\in\mathscr{L}$, and it satisfies the differential equation
\begin{equation}\wp'(w)^2=4\wp(w)^3-g_2(\omega_1, \omega_2)\wp(w)-g_3(\omega_1, \omega_2)\end{equation}
for \emph{invariants}
\begin{equation}
g_2(\omega_1,\omega_2)\stackrel{def.}{=}60\displaystyle\sum_{\omega\in\mathscr{L}-\{0\}} \frac{1}{\omega^4}, \ \ g_3(\omega_1, \omega_2)\stackrel{def.}{=}140 \displaystyle\sum_{\omega\in\mathscr{L}-\{0\}}\frac{1}{\omega^6}\end{equation}
where, moreover,
\begin{equation}g_2(\omega_1, \omega_2)^3-27 g_3(\omega_1, \omega_2)^2\neq 0.\end{equation}

Conversely, it is an amazing fact that if two complex numbers $g_2$ and $g_3$ are given that satisfy the condition $g_2^3-27g_3^2\neq 0$, then there exists a pair of non-zero complex numbers $\omega_1, \omega_2$ with $Im \ \omega_2/\omega_1>0$ such that $g_2(\omega_1, \omega_2)=g_2$ and $g_3(\omega_1, \omega_2)=g_3$, for $ g_2(\omega_1, \omega_2)$ and $ g_3(\omega_1, \omega_2)$ defined in (A.3) with respect to the lattice $\mathscr{L}=\mathscr{L}(\omega_1, \omega_2)$ generated by $\omega_1$ and $\omega_2$.  Thus from $g_2$ and $g_3$ one can also construct the corresponding $\wp-$function $\wp(w; \omega_1, \omega_2)$ (according to definition (A.1)), which in this case we also denote by $\wp(w; g_2, g_3)$ -- as we have so done in the previous sections.

Associated with $\wp(w)$ are the Weierstrass sigma and zeta functions $\sigma(w)$ and $\zeta(w)$, respectively:
\begin{eqnarray}
\zeta'(w)\stackrel{def.}{=}-\wp(w), && \displaystyle\lim_{w\rightarrow 0} \left( \zeta(w)-\frac{1}{w}\right)\stackrel{def.}{=}0,\notag\\
&&\label{eq: defnzetasigma}\\
\notag\frac{\sigma'(w)}{\sigma(w)}\stackrel{def.}{=}\zeta(w), && \displaystyle\lim_{w\rightarrow 0} \frac{\sigma(w)}{w}\stackrel{def.}{=}1.\end{eqnarray}


\end{document}